### v1

# Simultaneous elections make single-party sweeps more likely

Abstract
In a country with multiple elections, it may be expedient to hold some or all of them on a common polling date. Our main result, Theorem C, is that under certain assumptions, an increase in the simultaneity of polling increases the likelihood of a sweep, i.e. the likelihood that a single party wins all the elections. We discuss the applicability of our result to the two most common real world electoral systems, namely first-past-the-post (most voters, including US and India) and party list proportional representation (most countries).
We deduce Theorem C from a certain inequality proved in Theorem D, which is of independent interest. In particular, we connect our inequality to the Harris correlation inequality, which is a multivariate generalization of the Chebyshev sum inequality, and plays an important role in statistical mechanics and graph theory. More precisely, we show that Theorem D also implies Theorem F, which extends the domain of the Harris inequality to a larger class of functions.
MSC Classification: 60C05 60E15 91B12.  JEL Classification: D72
Keywords: Simultaneous elections, aligned functions, Harris inequality.
1 Introduction
Our main theme is that, under certain assumptions, a more simultaneous polling schedule makes it more likely that one party wins all the elections. We make this precise in three scenarios with increasing degree of generality.
1.1 Main results
Country A has two political parties and two national elections, Presidential and Parliamentary; in each election the candidate with the most votes wins. The election commission is considering whether to hold the elections simultaneously or on two separate polling dates. Each voter $h$ has a fixed probability $p_{h}$ of voting on any poll date [^1] This assumption is very much in the spirit of Banzhaf [4] and Penrose [30]., and the population is polarized in the sense that each voter prefers the same party for both elections. We define a sweep to be the outcome in which the same party wins both the presidency and a parliamentary majority.
Theorem A
The simultaneous election has a higher222throughout, “higher” is to be understood in the weak sense of “greater than or equal to.” likelihood of a sweep.
Country B has several political parties and several elections with varying notions of what constitutes a “win”. As before, each voter has a fixed probability of voting on any poll date and always votes for the same party. Moreover, the following two conditions hold for each election:
(a) for any voter turnout, at most one party can win;
(b) for any voter turnout, an additional vote for one party cannot turn its win into a non-win, nor turn some other party’s non-win into a win.
The election commission is exploring various polling schedules. We will say that schedule $\pi$ is coarser than schedule $\pi^{\prime}$ if any two elections that are simultaneous (have the same poll date) in $\pi^{\prime}$ are also simultaneous in $\pi$. We (re)-define a sweep to be the scenario in which one party wins all the elections.
Theorem B
A coarser polling schedule has a higher likelihood of a sweep.
Country C has a setup entirely similar to country B. However the election commission is considering staggered schedules in which each election might be spread out over several days, with different poll dates for different regions. We will say that staggered schedule $\Pi$ is coarser than $\Pi^{\prime}$ if any two elections that are simultaneous for a voter $h$ in $\Pi^{\prime}$ are also simultaneous for $h$ in $\Pi$.
Theorem C
A coarser staggered schedule has a higher likelihood of a sweep.
In a country with many political parties, it often happens that several parties come together in an alliance or coalition in order to form a government. Here one can distinguish two kinds of alliances – a pre-poll alliance which involves seat sharing, and a post-poll alliance in which a group of parties comes together after having fought the elections separately, perhaps compteing against each other for more seat. In this case, one may ask whether Theorem C continues to hold if we treat each coalition as single party? In section 5.6 we show that is indeed the case for many real -world electoral systems.
1.2 Structure of the paper
In section 5 we discuss the applicability of our results to various electoral systems, especially the US and Indian electoral systems, and contrast our results to other important themes in the literature such as social choice theory [3, 14].
As for the proofs, we note first that Theorem C implies Theorems A and B. In section 2 we reduce the proof of Theorem C to a mathematical result, Theorem D, which we then prove in section 3. Finally, in section 4, we discuss some further consequences of Theorem D, which might be of independent interest. This includes Theorem F which generalizes [^3] Theorem D generalizes the Harris inequality to correlations involving a larger class of functions. This is in contrast to the well-known FKG inequality [15] which extends the Harris inequality to a larger class of probability measures. One of the authors of the present paper has proposed a further generalization of the FKG inequality to several functions and obtained some partial results in this direction. However the general problem remains open, and seems quite hard; see [22, 33, 34, 35, 36] the Harris inequality – a well-known correlation inequality with applications in many different areas such as statistical mechanics, percolation theory, graph theory, and combinatorics [2, 17].
2 Proof of Theorem C
Let $H=\{1,\ldots,m\}$ and let $L=\{1,\ldots,n\}$ be the sets of voters and elections. The voter turnout, $T\subset H\times L,$ is the set of pairs $(h,l)$ such that voter $h$ has voted in election $l$. This turnout can be equivalently described by the tuple $(L_{1},\ldots,L_{m})$ where $L_{h}\subset L$ is the set of elections in which voter $h$ has voted, or by the tuple $(H_{1},\ldots,H_{n})$ where $H_{l}\subset H$ is the set of voters who have voted in election $l$.
Indeed, if $\mathcal{H},\;\mathcal{L}$ and $\mathcal{T}$ are the power sets (sets of all subsets) of $H,\;L$ and $H\times L$, then we have natural bijections $\lambda:\mathcal{L}^{m}\rightarrow\mathcal{T}$ and $\theta:\mathcal{H}^{n}\rightarrow\mathcal{T}$ given by
$\lambda(L_{1},\ldots,L_{m})=\bigcup\nolimits_{h}\{(h,l):l\in L_{h}\},\quad\theta(H_{1},\ldots,H_{n})=\bigcup\nolimits_{l}\{(h,l):h\in H_{l}\}.$ (1)
A staggered schedule determines a sequence $\Pi=(\pi_{1},\ldots,\pi_{m})$, where $\pi_{h}=\{M^{h}_{1},M^{h}_{2},\ldots\}$ is a partition of $L$ into subsets grouped by common poll dates for $h$. The probabilities $\mathbf{p}=(p_{1},\ldots,p_{m})$ induce a probability measure on $\mathcal{T}$ as follows.
Definition 1
We write $\mu_{\Pi,\mathbf{p}}(T)$ for the probability of obtaining $T$ by the following random procedure [^4] It is not too difficult to give an explicit product formula for $\mu_{\Pi,\mathbf{p}}(T)$ in terms of $\lambda^{-1}(T)=(L_{1},\ldots,L_{m})$, but we choose to skip this since it will not play any role in the proof.: For each $h\in H$ and for each $M^{h}_{i}$ in $\pi_{h}$, toss a coin with probability $p_{h}$ of heads, and if heads then include the set $\{(h,l):l\in M^{h}_{i}\}$ in $T$.
We define the “sweep” function for party $t$ on $\mathcal{T}$ as follows:
$X^{t}:T\xrightarrow{\theta^{-1}}(H_{1},\ldots,H_{n})\mapsto\chi_{1}^{t}(H_{1})\cdots\chi_{n}^{t}(H_{n}).$ (2)
where $\chi_{l}^{t}(H^{\prime})=1$ if party $t$ wins election $l$ when voters from $H^{\prime}$ cast their votes, and $\chi_{l}^{t}(H^{\prime})=0$ otherwise. Then turnout $T$ yields a sweep for party $t$ if and only if we have $X^{t}(T)=1$, and the probability of a sweep by party $t$ is the expectation
$\sum\nolimits_{T}X^{t}(T)\mu_{\Pi,\mathbf{p}}(T).$
We say a real function $f$ on $\mathcal{H}$ is increasing (resp. decreasing) at $h\in H$ if for all $S\subset H$ one has $f(S\cup\{h\})\geq f(S)$ (resp. $f(S\cup\{h\})\leq f(S)$). By our assumptions the functions $\chi^{t}_{1},\ldots,\chi^{t}_{n}$ are all increasing at $h$ if voter $h$ prefers party $t$, otherwise they are all decreasing at $h$. This motivates the following definition.
Definition 2
We say that the functions $f_{1},\ldots,f_{n}$ on $\mathcal{H}$ are aligned if for each $h\in H$, either they are all increasing at $h$ or all decreasing at $h$.
In the next section we will prove the following result.
Theorem D
Suppose $f_{1},\ldots,f_{n}$ are aligned and nonnegative on $\mathcal{H}$, and define
$F:T\xrightarrow{\theta^{-1}}(H_{1},\ldots,H_{n})\mapsto f_{1}(H_{1})\cdots f_{n}(H_{n}),\quad E(\Pi,\mathbf{p})=\sum\nolimits_{T}F(T)\mu_{\Pi,\mathbf{p}}(T).$
If $\;\Pi^{\prime}$ is coarser than $\Pi$ then we have $E(\Pi^{\prime},\mathbf{p})\geq E(\Pi,\mathbf{p})$.
Proof. (of Theorem C.) By Theorem D and the preceding discussion it follows that if we switch from $\Pi$ to $\Pi^{\prime}$, then this increases the probability of a sweep by party $t$ for every $t$. This implies Theorem C.
3 Proof of Theorem D
Proof. (of Theorem D.) It suffices to prove $E(\Pi^{\prime},\mathbf{p})\geq E(\Pi,\mathbf{p})$ for the special case where
  1. 1.
    $\Pi^{\prime}=\{\pi^{\prime}_{h}\}$ and $\Pi=\{\pi_{h}\}$ differ only for a single $h$, say $h=1$;
  2. 2.
    $\pi^{\prime}_{1}$ is obtained by combining two parts of $\pi_{1}$, say the first two, as follows:
    $\pi_{1}=\{M_{1},M_{2},M_{3},\ldots,M_{k}\},\quad\pi_{1}^{\prime}=\{M_{1}\cup M_{2},M_{3},\ldots,M_{k}\}.$
Indeed, by iterating 2) we obtain the inequality for any $\pi^{\prime}_{1}$ coarser than $\pi_{1}$. Now by iterating 1) we obtain the inequality for any $\Pi^{\prime}$ coarser that $\Pi$.
Thus we may assume that $\Pi$ and $\Pi^{\prime}$ are as in 1) and 2). Now the measures $\mu_{\Pi,\mathbf{p}}$ and $\mu_{\Pi^{\prime},\mathbf{p}}$ involve the same coin tosses for $\pi_{2},\ldots,\pi_{m}$ and for $M_{3},\ldots,M_{k}$ in $\pi_{1}$. It suffices to show that for each subset $T^{\prime}$ of $L\times H$ resulting from these tosses, the conditional expectation of $F(T)=f_{1}(H_{1})\cdots f_{n}(H_{n})$ is higher if we toss a single coin for $M_{1}\cup M_{2}$ rather than separate coins for $M_{1}$ and $M_{2}$.
To study this we write $\theta(T^{\prime})=(H_{1},\ldots,H_{n})$, and we set
$a_{i}=\prod_{l\in M_{i}}f_{l}(H_{l}\cup\{1\}),\;b_{i}=\prod_{l\in M_{i}}f_{l}(H_{l}),\text{ for }i=1,2;\quad c=\prod_{l\notin M_{1}\cup M_{2}}f_{l}(H_{l}).$
Then the conditional expectations for the single and double toss are, respectively,
$A=[pa_{1}a_{2}+(1-p)b_{1}b_{2}]c,\;B=[p^{2}a_{1}a_{2}+p(1-p)(a_{1}b_{2}+b_{1}a_{2})+(1-p)^{2}b_{1}b_{2}]c,$
where $p=p_{1}$. Now by an easy calculation we get
$A-B=p(1-p)(a_{1}-b_{1})(a_{2}-b_{2})c.$
Since $0\leq p\leq 1$, we have $p(1-p)\geq 0$, and since $f_{i}\geq 0$, we have $c\geq 0$. Further, since the $f_{i}$ are aligned, either they are all increasing or all decreasing, at $h=1$. In the first case we have $a_{i}\geq b_{i}$ and in the second case $a_{i}\leq b_{i}$, for $i=1,2$. In either case we get $(a_{1}-b_{1})(a_{2}-b_{2})\geq 0$, which implies $A-B\geq 0$.
4 Further consequences of Theorem D
For $n=2$, the set $L=\{1,2\}$ has only two partitions, namely $\{L\}$ and $\{\{1\},\{2\}\}$. In this setting, if $S$ is a subset of $H$ then we can define a partition sequence $\Pi_{S}=(\pi_{h})$ of $L$ as follows: $\pi_{h}=\{L\}$ if $h\in S$, and $\pi_{h}=\{\{1\},\{2\}\}$ if $h\notin S$.
The power set $\mathcal{T}$ of $H\times L$ can be identified with $\mathcal{H}\times\mathcal{H}$, and we write $\mu_{S}$ for the measure $\mu_{\Pi_{S},\mathbf{p}}$. Equivalently $\mu_{S}(H_{1},H_{2})$ is the probability of generating the pair $(H_{1},H_{2})$ by the following random procedure: For each $h\in S$ toss a coin with probability $p_{h}$ of heads, and if heads include $h$ in both $H_{1}$ and $H_{2}$, for $h\notin S$ toss the coin twice, once for inclusion in $H_{1}$ and once for $H_{2}$.
Definition 3
For functions $f_{1}$ and $f_{2}$ on $\mathcal{H}$ we define their convolution $f_{1}\ast f_{2}$ as follows: $f_{1}\ast f_{2}(S)=\sum_{\mathcal{H}\times\mathcal{H}}f_{1}(H_{1})f_{2}(H_{2})\mu_{S}(H_{1},H_{2})$.
The following theorem generalizes a result from [9] for increasing functions.
Theorem E
If $f_{1},f_{2}$ are aligned and nonnegative then $f_{1}\ast f_{2}$ is increasing.
Proof. In the notation of Theorem D for $n=2$ we have $f_{1}\ast f_{2}(S)=E(\Pi_{S},\mathbf{p})$. Evidently $\Pi_{S^{\prime}}$ is coarser than $\Pi_{S}$ if and only if $S^{\prime}$ contains $S$. Thus $f_{1}\ast f_{2}$ is an increasing function.
Let $\mu$ be the product measure on $\mathcal{H}$ defined by
$\mu(S)=\prod_{h\in S}(p_{h})\prod_{h\notin S}(1-p_{h})$
and let $E(f,\mu)=\sum_{S}f(S)\mu(S)$ be the corresponding expected value. Then the following result generalizes the Harris inequality [17] for increasing functions.
Theorem F
If $f_{1},f_{2}$ are aligned and nonnegative then we have
$E(f_{1}f_{2},\mu)\geq E(f_{1},\mu)E(f_{2},\mu).$
Proof. A straightforward calculation shows that we have $f_{1}\ast f_{2}(H)=E(f_{1}f_{2},\mu)$ and $f_{1}\ast f_{2}(\emptyset)=E(f_{1},\mu)E(f_{1},\mu)$. Thus the result follows from Theorem E.
5 Discussion
5.1 Random Voter Turnout
A key hypothesis of our model is that the probabilites of turning out to vote are independent across individuals (though they may vary arbitrarily across them). This is very much in the spirit of Penrose [30], and later Banzhaf [4]), both of whom considered the special case where these probabilities were identically $1/2$
The recent “instrumental theory of turnout” [^5] The theory posits individuals for whom the act of voting is “instrumental” to maximizing their own “payoffs”, though these payoffs may be very broadly defined, incorporating not only the joy of seeing their party win but also the joy of participating in the election (alongside the cost of participating). provides theoretical underpinnings for the independence assumption. Some of the analyses in the field are focused on individual rational choice (in an exogenously fixed environment), modeled in terms of adaptive learning (e.g., [19]), or utility maximization (e.g., [39]) or minmax regret (e.g.,[12]). Others discuss full-blown strategic interaction among the voters (and sometimes also candidates) and derive Nash equilibria of “participation games” (e.g., [21], [26], [27]). These analyses differ considerably from one another on how individual turnout probabilities are formed, but they all imply the independence of those probabilities. (See the surveys in [8] and, for a broader panorama, in [16]). Some recent models (e.g., [31]), while maintaining the instrumentality assumption, introduce the possibility of communication — broadly defined — between candidates, media and voters. This gives rise to correlated equilibria where the turnout probabilities of voters are no longer independent. (Also, there are several behavioral models which openly depart from the “instrumental theory” and directly incorporate correlated turnout among the voters, such as voting in “teams” (see [11], [37] and the references therein).
The independence hypothesis is crucial to our analysis. However there is a “special kind” of correlation that can be admitted in our model (which, while mathematically trivial, may be useful conceptually and in applications). As in Harsanyi ([18]) define the type of voter $h\in H$ to consist of two components: the party $t\in T$ that $h$ will vote for; and the probability $p_{h}$ with which $h$ will turn out to vote on any polling date (across all elections). Prior to the announcement of the polling schedule, let nature pick the vector of voter-types from some apriori fixed set according to a probability distribution. [^6] Two interpretations come to mind. (I) Think of nature’s “moves” as public announcement of news which affects voter behavior, increasing or decreasing their turnout probabilities and sometimes even determining which party they will vote for. (II) Suppose there is full correlation in turnout probabilities, indeed — as an extreme case — suppose even that some set of voters go to the polls with probability $1,$ while the others go with probability 00 (i.e., abstain). Our model does admit this scenario, though now the likelihood of a sweep (by any party) will be invariant of the polling schedule. However, were nature to add idiosyncratic (independent) uncertainty — no matter how small — in voters’ turnout behavior, then simultaneous polls would induce strictly higher likelihood of a sweep compared to any other polling schedule. Indeed it would suffice to add uncertainty for at least one voter $h$ who is not a “dummy” in the sense that he affects the electoral outcome with positive probability, i.e., $\chi_{l}^{t}(Z\cup\left\{h\right\})-\chi_{l}^{t}\left(Z\right)\neq 0$ for some $t,l$ and some turnout $Z$ which occurs with positive probability. For then the probability of a sweep by party $t$ would be higher under simultaneous polls than under any other schedule. Thus nature’s move is ex ante: every voter’s type stays fixed throughout the elections. Now, if schedule $\Pi$ is coarser than $\Pi^{\prime}$, then the probability of a sweep is higher under $\Pi$ than under $\Pi^{\prime}$ contingent on every move of nature (by Theorem C), therefore obviously also in expectation across all the moves of nature. Note that by allowing the type of a voter to vary, we have made some room for undecided voters who could vote for different parties. However if nature were to move ex post, repeatedly and independently between different poll dates, the argument given above breaks down.
5.2 Electoral systems
There is a wide variety of election methods in use across the world. For an extensive survey, covering all the methods we cite below, see [1].
Our analysis applies to first-past-the-post elections that are held in several countries including India, USA, Canada and UK (the first two of which we shall discuss below in some detail).
It also applies to elections based on proportional representation, where each person votes for a single party of his choice, and the percentage of parliamentary seats accorded to any party equals (as nearly as possible) the percentage of total votes it received. Of course, for our analysis to be valid, the method for “rounding” the seats (into integers) should be “monotonic”  [^7] Monotonicity holds for most of the rounding methods used in practice, e.g., the Highest Average or the Largest Remainder Methods [1] in the obvious sense, so that condition (b) is not violated; and, furthermore, condition (a) must also hold (e.g., the party with the most seats is deemed to win). Such elections are in use — with minor modifications — in many countries and also in the European Parliament (see, again, [1] for details).
There is, however, a category of elections that lies outside the ambit of our model, and it includes the Two Round System (used in the French Presidential Election), Single Transferable Vote, Alternative Vote, Supplementary Vote, Borda Count and so on (described in detail in [1]). In all these cases, a vote is tantamount — either implicitly or explicitly — to a (possibly partial) ranking of the candidates/parties involved. Such elections have received the lion’s share of attention in the theoretical literature on “social choice” (SC) ever since Arrow’s pioneering monograph [3]. (For a succinct survey, see, e.g., [14].) The concerns of the SC literature are quite different from ours, but even from the “high level” viewpoint, there are salient differences. First, our voter only specifies her/his top choice instead of ranking the candidates. More importantly, in SC, the entire voter population is assumed to always turn out to vote. In sharp contrast, it is the variable, stochastic turnout of voters on which our entire analysis turns. It might be interesting to examine, along the lines of this paper, how the probability distribution of electoral outcomes depends on the stochastic turnout of voters, in the context of the more sophisticated elections considered in SC (e.g., Borda Count, see [23]), but that would be a topic for another paper.
5.3 Single-Party Sweeps
A single-party sweep (also known as “one-party monopoly” or “one-party dominance”) is a widely prevalent phenomenon, both in local and national elections ([28], [29], [38]) Its causes have been discussed from different standpoints, including: the strategic timing of elections ( [20]), economic intervention ([20]), manipulation of electoral rules ([24] ), and fragmented opposition ([40]). There does not seem to be much in the literature connecting polling schedules to single-party sweeps, which is the focus of this paper. Two notable exceptions are the “coattail effect” in the context of US elections (see, e.g., [6, 7, 25]). and a recent study ([5]) of the Indian elections. In both these cases, the analysis turns (see the next two sections) on the postulate that a voter’s behavior changes when elections are held simultaneously, on account of either the popularity of the Presidential candidate or the “salience” of one of the political parties. In sharp contrast, voters’ behavior is fixed in our model, regardless of the polling schedule or the characteristics of the candidates standing for election. It is the statistical aggregate of voters’ behavior that nevertheless varies with the schedule and drives our result. Thus our analysis brings to light a cause for single-party sweeps that is not only novel, but invariant of the candidates’ influence on voters and — to that extent — more constant.
5.4 US Elections and The Coattail Effect
The data from USA shows that a single party is more likely to win multiple elections when they are held simultaneously. This phenomenon is referred to as the “coattail” effect, based on the the rationale that a popular presidential candidate fosters additional wins for his party in Congress, with many members of Congress voted into office “on the coattails” of the president. Indeed, during Congressional elections in USA that coincide with the presidential race, the party of a popular president tends to win more seats in Congress compared to elections held in the midterm of the Presidency (the latest instance of this being the US elections in November 2024).
Our analysis brings to light a completely different mechanism for a single-party sweep, which is not at odds with the coattail effect, but independent of it. Two behavioral causes are commonly postulated for the coattail effect: a popular presidential candidate can mobilize each member in his party base to turn out to vote with higher probability; and, moreover, he can sway undecided voters to his party’s fold. [^8] A cause of a different nature is the “anti-incumbency factor” so often alluded to by political analysts. It can come into play during midterm elections, especially if there is widespread perception that the ruling party has been unable to fulfil its electoral promises. For then the rival party makes gains in the midterm elections, further accentuating the benefit of simultaneous elections for the ruling party. These causes are ruled out in our model since the party that $h$ will vote for, as well the probability $p_{h}$ that he will vote on any poll date, are both exogenously fixed for every $h\in H$ and not susceptible to change. It is purely “the statistics of voting” that causes the single-party sweep. Were the coattail effect to be present, it would further boost the probability of a sweep for the favored party [^9] We could incorporate the “coattail” in our model as an increase in the probabilities of turnout of the adherents of the favored party in the Presidential election and in (at least) all other contemporaneous Congressional elections (while the turnout probabilities of adherents of rival parties stay fixed or go down). This will obviously increase the probability of win for the favored party in every election. In this sense, our model is not opposed to the coattail effect, but accomodative of it.; but even in its absence, the likelihood of a sweep increases as the elections are made “more simultaneous” in the precise sense of Theorems B and C.
The probability of a single-party sweep can, of course, be very small if there are many parties or many elections. It becomes significant when, as is often the case, there are two parties contending two (or three) elections. To the extent that “Duverger’s Law” operates and a representative democracy settles into a two-party system, the sweep can be quite a robust phenomenon [10, 13, 32].
5.5 Indian Elections: “ONOE” & “Double-Engine” Governance
Elections in India are held as follows. Political parties contend for seats in the Lok Sabha at the national level and also in $28$ Vidhan Sabhas, one for each state. The country (resp., state) is partitioned into electoral districts for the Lok (resp., Vidhan) Sabha election, the candidate with the most votes is elected from each district, and the party with the most candidates is deemed to have won the election. [^10] W.l.o.g. we may assume that the entire electorate of India participates in each Vidhan Sabha election, viewing voters outside the state as “strategic dummies” whose votes do not count. (Condition (b) still holds for such dummies, and therefore our model applies.)
The possibility of simultaneous elections is no longer theoretical but it looms large in India and is the subject of a fierce ongoing debate, set into motion by the BJP and its allies (currently in power in the Lok Sabha) seeking a Constitutional Amendment to implement their “One Nation, One Election” (ONOE) proposal (which would make all Vidhan Sabha elections simultaneous with the Lok Sabha election). At the same time, the BJP has been strongly campaigning for “double-engine” governance, especially during elections in the states. [^11] This expression signifies the scenario in which the same party governs in both the Center and the State (for then the two can presumably work more in concert — as a “double engine” pulling in the same direction — for the development of the State). Our Theorem C shows that the likelihood of one party winning the Lok Sabha and the Vidhan Sabha of any given state is higher when both elections are held simultaneously compared to when they are not. It immediately follows that, under ONOE, the probability of “double-engine” governance would go up for every state. This would be a systemic change in Indian elections, brought on by ONOE, for it would pertain no matter which party wins the Lok Sabha.
The study in [5] refers to the pronounced correlation in the Indian data between simultaneous elections and single-party sweeps (over those elections). To explain it, the authors posit a model of behaviorally-constrained voters for whom information acquisition is costly. If a party is salient in the Lok Sabha elections, then the electorate has a higher probability to also vote for it in any contemporaneous Vidhan Sabha election, rather than acquire costly information about its rivals which could alter that decision. Such is not the case if the Vidhan Sabha election is held at a very different time, for then the “salience factor” is missing and information acquisition becomes less costly. By fixing the type of a player during elections, we have ruled out this kind of varition in voter behavior. Our model highlights an alternative, purely stastistical cause of the correlation.
Finally recall that the notion of “win” is very general in our model and can vary across elections, subject only to conditions (a) and (b). This enables us to consider Indian elections from diverse viewpoints, including the following two starkly different ones: $1+28=29$ “literal” elections, or $2$ “figurative” elections (with the $28$ Vidhan Sabha elections viewed as one conglomorate). The literal elections clearly fit into our model, but here the likelihood of a sweep is so miniscule that it is unlikely for either party to be focused on it. Sweeps occur with higher ptobability in the $2$-election framework, where “win” in Vidhan Sabhas can be defined in diverse ways, e.g., (i) winning a specific set of states [^12] The contending parties may have different sets of states in their sights but, so long as the intersection of their sets is non-empty, condition (a) holds and our analysis remains intact.; (ii) winning most states overall; (iii) winning most seats overall (relevant for the future Rajya Sabha); and so on. All these scenarios are covered by Theorem C, which allows for staggered polling schedules (i.e., personal partitions of the elections in $L)$.
5.6 Party Alliances
Consider a partition of parties into sets, each of which constitutes an electoral alliance. Define the type of an alliance to be pre-poll if the parties in the alliance jointly field one candidate in every election, and do not contend against one another; and to be post-poll if they contend every election on their own, but come together afterwards by pooling their elected candidates. Our analysis will remain intact if conditions (a) and (b) hold, viewing each alliance as a single contender in all the elections. This is easily seen to be so in elections based on first-past-the post criterion or on proportional representation, provided the type of each alliance is fixed across all polling schedules (i.e., either it is pre-poll for all schedules, or post-poll for all). To check (b) in the presence of post-poll alliances (the only scenario where it is not immediately obvious), note that if a vote is cast in favor of an alliance it does not benefit any of the rival alliances, and therefore cannot hurt the alliance either, since the total number of elected candidates is fixed in every election.
However, if the set of parties that form an alliance, or the type of an alliance, varies across elections, or across the polling schedules, or with the electoral turnout, then our analysis breaks down.